# quality of service 5G fifth generation

### v1

@@ IV-C2 Contrastive Loss
Contrastive loss is a widely-used loss function in ML for various representation learning tasks, including image or text classification. Its fundamental aim is to construct a representation of data points that encourages similar data points to be closer together in the representation space while simultaneously pushing dissimilar data points farther apart [17].
Mathematically, the contrastive loss can be expressed as follows:
@@
To use contrastive loss in our resource allocation problem, we need to create similar and dissimilar channel gain matrices. One way to create channel gain matrices that are either similar or dissimilar is by using a statistical model of the channel that captures the relevant properties of the communication environment. The Rayleigh fading model is a commonly used statistical model for wireless channels, which assumes that the channel gain is a complex Gaussian random variable with zero mean and a variance that depends on the distance between the transmitter and receiver [47, 48].
To create similar channel gain matrices, coefficients with similar magnitude and phase can be generated. Conversely, to create dissimilar channel gain matrices, coefficients with different magnitude and phase can be generated by introducing randomness into the channel model by using a Rayleigh fading model. To generate a channel gain matrix using this model, a matrix of complex Gaussian random variables with zero mean and a variance that depends on the distance between the transmitter and receiver can be generated as well as considering the interferance as noise [49, 50]. The variance can be adjusted to create similar or dissimilar channel gain matrices based on the desired level of correlation between the channels. The mathematical formulation of this approach can be expressed as:
$\mathbf{H}_{noise}=\mathbf{H}+\mathbf{\epsilon},$ (37)
where $\mathbf{H}_{noise}$ is the channel gain matrix with added noise, $\mathbf{H}$ is the original channel gain matrix, and $\mathbf{\epsilon}$ is a Gaussian noise matrix with zero mean and standard deviation $\sigma$.
Once we have created similar and dissimilar channel gain matrices, we can train our contrastive loss function using these matrices as inputs. The loss function can be formulated as:
@@
where $\theta$ is the set of parameters in the variational autoencoder model, $\tau$ is a temperature hyperparameter and $L(\theta)$ is the contrastive loss function. $\underline{H_{i}}$ and $\bar{H_{i}}$ are the similar matrices to the the main channel gain matrix ($H_{i}$), which allow us to train our proposed model [51]. By minimizing this loss function, we can learn a set of parameters that enable our model to differentiate between similar and dissimilar channel gain matrices and produce a power control decision that optimizes the network performance.
IV-C3 Training the proposed model
In this section, we will describe the training phase of our proposed method, which consists of two parts: the supervised learning part, where the VAE is trained to regress the power transmission, and the unsupervised learning part, where the contrastive loss is used to improve the generalization and robustness of our proposed method. There are several key parameters that should be taken into consideration in training the supervised part. The latent space dimension represents the size of the hidden representation learned by the VAE [45]. A larger latent space dimension can lead to a more expressive model, but also increases the risk of overfitting. The learning rate determines the speed at which the model learns from the data. The batch size determines the number of samples used for each iteration of the training process. The number of epochs determines the number of times the model will be trained on the data. $L1$ and $L2$ regularization to prevent overfitting by adding a penalty term to the loss function of supervised learning [52]. The proposed model is developed using the PyTorch [53] software framework. Both validation size and test size are $20\%$ of the dataset, allowing for a more thorough evaluation of the proposed model’s performance. Hyper-parameter tuning yields the value of hyperparameters, which can be found in the Table III.
The initial model was pre-trained through self-supervised learning, using unlabeled channel gain matrices that corresponded to the stated loss function. The pre-training was based on regression loss. After pre-training, the DL model was fine-tuned for power transmission regression using the proposed method. The performance of the trained model was evaluated on test data points using various performance metrics, while the number of labeled training channel gain matrices were varied. Additionally, a supervised baseline performance for the regression task was obtained by training the DL model with all available labeled training channel gain matrices.
@@ IV-C4 Performance Metrics
The performance metrics for evaluating our proposed method are Mean Absolute Error (MAE), cosine similarity, and correlation. MAE measures the average absolute difference between the predicted values and the actual values. Cosine similarity measures the similarity between two vectors in a multi-dimensional space. The Pearson correlation ($R$) represents the amount of variation in the output-dependent attribute that can be predicted by the input-independent variable. It also provides insight into how accurately a model can predict outcomes based on how much of the variation in those outcomes can be explained by the input-independent variable. They can be calculated as follows:
$MAE=\frac{1}{N}\sum_{i=1}^{N}|y_{i}-\hat{y}_{i}|,$ (39)
@@
$correlation(\hat{Y},Y)=\frac{\text{cov}(\hat{Y},Y)}{\sigma_{\hat{Y}}\sigma_{Y}},$ (41)
where $N$ is the total number of samples, $y_{i}$ is the true value, and $\hat{y}_{i}$ is the predicted value, and $\cdot$ represents the dot product, and ||$\cdot$|| ⋅ | represents the magnitude of a vector. As for correlation, $\text{cov}(\hat{Y},Y)$ is the covariance between the predicted value and the true value, and $\sigma_{\hat{Y}}$ and $\sigma_{Y}$ are the standard deviations of the variables.
@@ V SIMULATION RESULTS
In the following subsections, we will begin by analyzing the initial values to simulate the proposed algorithm. Subsequently, we will compare the results obtained using SS-VAE with those from ESA and DQN. This comparative analysis aims to assess the accuracy and effectiveness of our proposed DL model in addressing the resource allocation problem.
@@ V-A Initial points and performance of training the model
In this section, we present numerical results for the considered resource allocation problem to evaluate the performance of our proposed model (SS-VAE) compared to state-of-the-art methods, namely ESA and DQN. The initial values and parameters used to run and simulate our proposed algorithm (SS-VAE) and the benchmark algorithms (ESA and DQN) are provided in Table III, Table III, and Table III.
TABLE I: Resource Allocation Simulation Parameters
TABLE II: DQN Training Parameters
TABLE III: Proposed Model Training Parameters
@@
TABLE II: DQN Training Parameters
TABLE III: Proposed Model Training Parameters
As previously discussed, our proposed model aims to estimate the transmission power corresponding to each input channel matrix. These estimated power values, derived from the weights of the VAE, are then used to predict the allocation indicators ($\alpha_{u,b,s},\beta_{u,b,m,s}$). To streamline the training process, it is crucial to convert the dimensions of each channel matrix data into a one-dimensional vector. More specifically, our model takes as input a flattened $UB$-dimensional vector obtained from the channel gain matrix.
The loss values for both the validation and training data points is depicted in Figure 4.
@@
The decreasing trend of loss values illustrate that the proposed method (SS-VAE) could accurately estimate the target values (power of transmission, association indicators). In fact, the model can map the input channel gain matrix to the power of transmission as well as the indicators by employing the VAE’s property in learning the underlying information.
The performance metrics for the test data, as presented in Table IV, clearly indicate that our proposed method (SS-VAE), which estimates transmission power based on the input channel gain matrix, exhibits the best generalization. This improvement can be attributed to the utilization of VAE, which has been proven to enhance generalization in various related tasks. In addition, the pre-training method involving the contrastive loss has further enhanced our model’s capacity to learn the underlying features, ultimately reducing generalization errors.
TABLE IV: Performance metrics of the proposed model
@@
The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions. This suggests that a substantial proportion of the variability in the dependent variable (power of transmissions) can be accounted for by the independent variables (the channel gain matrix) integrated into the model. In simpler terms, the model is effective at describing and understanding the variations in the data, especially in terms of transmission power.
@@ V-B Comparison of the algorithms
In the next phase, we will assess the performance of our proposed SS-VAE method by comparing it to the ESA algorithm. This evaluation will specifically consider the impact of the number of labeled samples used for training the supervised part of SS-VAE, as illustrated in Figure 5.
@@
We analyze the aggregated throughput under varying numbers of labeled data points, as generated by our proposed SS-VAE algorithm and the ESA. As depicted in Figure 5a, our proposed model achieves a similar aggregated throughput when compared to the ESA. Moreover, the discrepancy between our model and the ESA decreases significantly as the number of labeled data points increases.
We further evaluate the performance of the SS-VAE model and ESA in terms of binary variables related to user association ($\alpha_{u,b,s}$) and PRB allocation ($\beta_{u,b,m,s}$). Figure 5b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them ($\frac{|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|}{\text{\# Number of }(\alpha\&\beta)}$), as the number of labeled channel gain matrices (data points) increases. Notably, the difference between the $L1-norm$ of binary variables obtained by the SS-VAE method significantly decreases in comparison to ESA as the number of labeled data points grows.
The results above highlight the reliable performance of our proposed method (SS-VAE) in estimating transmission powers, user association, and PRB allocation when compared to the ESA. In the subsequent section, we will assess its performance under different initial conditions, varying the number of slices, UEs, and thresholds for transmission power and rate. Since the SS-VAE exhibits lower error compared to the ESA but doesn’t perform identically in the resource allocation problem, we will also evaluate its performance alongside another benchmark algorithm, the aforementioned DQN algorithm in Section IV-B. DQN has demonstrated practical and valuable results in resource allocation problems within wireless networks and similar domains, enabling us to assess the robustness of our proposed algorithm.
@@ V-C Performance results
In this section, we will assess the performance of the proposed model (SS-VAE) in the resource allocation problem under various scenarios. To begin, we will examine the aggregate weighted rate in relation to the number of UEs using three different algorithms: SS-VAE, ESA, and DQN, as depicted in Figure 6.
@@
Figure 6a demonstrates that the proposed approach (SS-VAE) exhibits a performance that is comparable to the ESA. There are no significant differences in terms of aggregated throughput between the two. Additionally, SS-VAE achieves a superior aggregated throughput across all UE numbers when compared to the DQN algorithm. This highlights the robustness and high accuracy of our proposed method for the resource allocation problem. It’s worth noting that the sum of data rate increases initially as more UEs join the service for all three examined algorithms. However, it reaches a plateau after 30-35 UEs per service due to power limitations and interference constraints.
Figure 6b depicts the aggregated throughput achieved by the SS-VAE algorithm concerning different power levels of O-RU and varying numbers of UEs. As per established knowledge [15], both increasing the power of O-RU and the number of UEs lead to higher aggregated throughput. The highest aggregated throughput is observed when the power of O-RU is at its highest.
Figure 7 displays the aggregated throughput for two specified service types (eMBB and URLLC), which has been obtained using the SS-VAE and DQN algorithms, in relation to the power values. This presentation aims to evaluate the performance of our proposed algorithm for these services in comparison to a benchmark algorithm.
Figure 7 displays the aggregated throughput for the two considered service types, namely eMBB and URLLC, as determined by the SS-VAE and DQN algorithms. This visualization is intended to evaluate the performance of our proposed algorithm (SS-VAE) for these services in comparison to a benchmark algorithm.
@@
Figure 7a depicts the influence of O-RU’s maximum power thresholds on the aggregate weighted rate for both eMBB and URLLC services, comparing two algorithms, SS-VAE and DQN. The results indicate a direct correlation between increased maximum power and higher weighted rates. Additionally, the SS-VAE algorithm consistently outperforms the DQN algorithm, achieving higher aggregated throughput for both service types across various power values.
Similarly, in Figure 7b, variations in total data rates are displayed concerning the maximum power for distinct cases of eMBB and URLLC services. The results reveal a direct increase in data rates as the maximum power of each UE in different service instances rises. Notably, the SS-VAE algorithm consistently outperforms the DQN algorithm, securing higher aggregated throughput for both service types across various power threshold values.
VI CONCLUSION
@@
In this paper, a novel model is introduced for the O-RAN system’s downlink, aiming to coordinate multiple independent xAPPs to tackle the resource allocation problem with network slicing in the O-RAN architecture. This problem is addressed for two specific 5G services, namely eMBB and URLLC, with the overarching objective of optimizing the aggregate weighted rate by managing RU associations, power allocation, and PRBs.
@@
Due to the non-convex and NP-hard nature of the resource allocation optimization problem, we introduced a semi-supervised learning approach as an efficient solution for UE and PRB allocation. Our two-step algorithm starts by reformulating the problem into an optimization problem considering the objective function and constraints. The second step consists of a two-part training process: firstly, supervised learning employing a VAE to estimate power transmission regression and association indicators, and then unsupervised learning with a contrastive loss approach, enhancing the model’s generalization and robustness.
To evaluate our proposed algorithm’s performance, we conducted comparisons with benchmark algorithms ESA and DQN across various scenarios. The results indicate that our approach achieves a rate similar to ESA while outperforming the DQN algorithm, highlighting its effectiveness. We further explore different scenarios tailored to each service, eMBB and URLLC, considering their unique QoS requirements. Through simulations, we demonstrate the superior efficiency of our model, particularly in terms of sample efficiency, making it a compelling solution for resource allocation challenges within the O-RAN system.
@@
In future research, we plan to explore several areas. Firstly, we intend to evaluate the end-to-end delay experienced by individual UEs and investigate the impact of deployment costs associated with network function placement tackle. Furthermore, addressing UE mobility is crucial in practical scenarios. In addition, we will explore alternative ML and optimization algorithms to improve our resource allocation approach, providing new research avenues in this dynamic field.

### v2

@@ IV-C2 Contrastive Loss
Contrastive loss is a widely-used loss function in DL for various representation learning tasks, including image or text classification. Its fundamental aim is to construct a representation of data points that provokes similar data points to be closer together in the representation space while simultaneously pushing dissimilar data points farther apart [17]. Mathematically, the contrastive loss can be expressed as follows:
@@
To apply contrastive loss in our resource allocation problem, we need to create similar and dissimilar channel gain matrices. This can be achieved using a channel’s statistical model that reflects the relevant properties of the communication environment. By introducing randomness through a Rayleigh fading model, we can produce channel gain matrices with varying coefficients in magnitude and phase, thus creating either similar or dissimilar matrices. To generate a channel gain matrix using this approach, we create a matrix of complex Gaussian random variables with zero mean, where the variance is influenced by the distance between the transmitter and receiver while also accounting for interference as noise [45, 46]. The variance can be adjusted to control the level of correlation between the channels, thereby generating similar or dissimilar channel gain matrices as needed. We can train our model with contrastive loss function once we have created similar and dissimilar channel gain matrices. The loss function can be formulated as:
@@
where $\theta$ is the set of parameters in the VAE model, $\tau$ is a temperature hyperparameter, and $L(\theta)$ is the contrastive loss function. $\underline{H_{i}}$ and $\bar{H_{i}}$ are similar matrices to the main channel gain matrix ($H_{i}$) and are used in training the SS-VAE model [47]. By minimizing this loss function, we can find a set of parameters that enable our model to differentiate between similar and dissimilar channel gain matrices; therefore, it is used to solve the resource allocation problem.
The connection between the VAE and the additional contrastive loss is essential for an optimal solution. The initial step establishes a strong base by using supervised learning to acquire a good representation of the resource allocation problem. The subsequent step further improves these representations, enhancing the model’s capacity to generalize and effectively handle new, unforeseen scenarios.
IV-C3 Training the SS-VAE model
The training phase of SS-VAE method consists of two parts. First, supervised learning involves training the VAE and using the VAE’s Encoder to predict power transmission. The second part is unsupervised learning, which uses contrastive loss to improve the generalization and robustness of our method.
The latent space dimension represents the size of the hidden representation learned by the VAE [48]. A larger latent space dimension can lead to a more expressive model but also increases the risk of overfitting. The learning rate determines the speed at which the model learns from the data. The batch size determines the number of samples used for each iteration of the training process. The number of epochs determines the number of times the model will be trained on the data. $L1$ and $L2$ regularization to prevent overfitting by adding a penalty term to the loss function of supervised learning [49]. The model is developed using the PyTorch [50] software framework. The dataset’s validation and test sizes are $20\%$, allowing for a more thorough evaluation of the model’s performance. Hyperparameter tuning yields the value of hyperparameters, which can be found in Table IV.
@@ V SIMULATION RESULTS
We analyzed the initial values in the following subsections to simulate the SS-VAE algorithm. Subsequently, we will compare the results obtained using SS-VAE with those from ESA and DQN.
@@ V-A Initial points and performance of training the model
This section presents the numerical results for the resource allocation problem to evaluate the performance of the SS-VAE model compared to ESA and DQN. In our training and testing scenarios, each RU in the O-RAN deployment is equipped with 25 PRBs and operates with a single antenna. We consider a network with two slices, eMBB and URLLC services. These UEs are randomly distributed within a single-cell environment, mostly 200 to 500 meters from the RU. The initial values and parameters used to run and simulate SS-VAE and the benchmark algorithms (ESA and DQN) are provided in Table IV, Table IV, and Table IV.
TABLE II: Resource Allocation Simulation Parameters
@@
Figure 4 shows the training and validation loss values. The decreasing trend of loss values indicates that the SS-VAE method could accurately estimate the target values (power of transmission, association indicators). The model can map the input channel gain matrix to the power of transmission and the indicators by employing the encoder part of the VAE’s to learn the underlying information.
The performance metrics for the test data, as given in Table V, clearly indicate that SS-VAE method obtained optimal performance in the test dataset, demonstrating its generalization ability. The $R^{2}_{Score}$ value for the test data indicates that our model successfully explains a significant portion of the variance in estimating the power of transmissions. This improvement can be attributed to the utilization of VAE, which has been proven to enhance generalization in various related tasks. Furthermore, the pre-training method involving contrastive loss has improved our model’s ability to learn the underlying features, reducing generalization errors.
TABLE V: Performance metrics of the SS-VAE model
@@ V-B Comparison of the algorithms
We evaluated the performance of our SS-VAE method compared to the ESA algorithm, focusing on how the number of labeled samples used for training affects the results, as shown in Figure 5. We analyze the aggregated throughput under varying numbers of labeled data points obtained by SS-VAE algorithm and the ESA. As depicted in Figure 5a, SS-VAE achieves a similar aggregated throughput compared to the ESA. Moreover, the discrepancy between SS-VAE and the ESA decreases significantly as the number of labeled data points increases. The difference in aggregated throughput between these methods is $2.1$ for $2000$ labeled samples, which will be decreased using more labeled samples.
@@
We also evaluate the performance of the SS-VAE and ESA based on binary variables related to user association ($\alpha_{u,b,s}$) and PRB allocation ($\beta_{u,b,m,s}$). Figure 5b displays the ratio of the absolute sum of differences in these binary variables between the SS-VAE and ESA to the total number of them ($\frac{|\alpha_{ESA}-\alpha_{SS-VAE}|+|\beta_{ESA}-\beta_{SS-VAE}|}{\text{\# Number of }(\alpha\&\beta)}$), as the number of labeled samples increases. The difference between them reaches its minimum at 2000 samples (4.7% error) and is expected to decrease significantly as more labeled samples are used.
These results demonstrate the reliable performance of our SS-VAE method in addressing the resource allocation problem. The following section will evaluate its performance under varying initial conditions, including different slice numbers, UEs, and transmission power and rate thresholds. Although SS-VAE shows lower error than ESA, it does not perform identically in resource allocation. Therefore, we will also compare it with the DQN algorithm, which has shown effective results in similar resource allocation problems within wireless networks [31, 35, 51].
@@ V-C Performance results
Figure 6 illustrates the aggregate weighted rate across different numbers of UEs as achieved by three algorithms: SS-VAE, ESA, and DQN. Figure 6a demonstrates the aggregated throughput obtained by SS-VAE, ESA, and DQN for different number of UEs. The difference in performance between SS-VAE and ESA is consistent across all UEs. Moreover, SS-VAE outperforms DQN in aggregated throughput across all UE. In addition, the aggregated throughput increases initially as more UEs join the service for all three examined algorithms. However, it reaches a plateau after 30-35 UEs per service due to power limitations and interference constraints. Figure 6b shows the aggregated throughput achieved by the SS-VAE algorithm at different maximum power levels of O-RU and different numbers of UEs. In resonation with [15], increasing the maximum power of O-RU and the number of UEs leads to higher aggregated throughput. The highest aggregated throughput is observed when the maximum power of O-RU is at its highest.
@@
Figure 7 displays the aggregated throughput for two specified service types (eMBB and URLLC), which has been obtained using the SS-VAE and DQN algorithms, about the maximum power of slice and RU. Figure 7a and Figure 7b both illustrate the aggregate weighted throughput achieved by SS-VAE and DQN under differetn maximum power levels for RU and slices, respectively, in eMBB and URLLC services. Both figures demonstrate a direct correlation between increased maximum power levels and higher weighted throughput. Across all scenarios, the SS-VAE algorithm consistently outperforms the DQN algorithm, providing higher aggregate throughput for both service types, regardless of whether the power levels pertain to the RU or the individual slices. Due to their need for higher data rates, eMBB service type exhibits higher aggregate throughput than URLLC services. This consistent performance advantage underscores the effectiveness of the SS-VAE algorithm in optimizing resource allocation across different power settings.
Although ESA ensures a globally optimal solution, its computational complexity makes it impractical for real-world applications. DQN, while powerful, faces challenges related to convergence time, exploration efficiency, and sample requirements. Our SS-VAE method employs the strengths of supervised and unsupervised learning to provide a scalable and efficient solution for resource allocation in O-RAN, outperforming these benchmarks in real-world scenarios. This contrasts to DQN, which may struggle with generalization due to its reliance on extensive exploration and sample efficiency. Additionally, DQN often requires significant training time to converge to an optimal solution, especially in environments with high variability and complex constraints. It may also face challenges with exploration efficiency, often needing many iterations to fully explore the solution space, resulting in suboptimal performance in dynamic conditions.
VI COMPUTATIONAL COMPLEXITY ANALYSIS
This section analyzes the computational costs of the ESA, DQN, and the SS-VAE method and discusses how each algorithm scales with increasing network size and complexity.
VI-1 ESA
The ESA evaluates all possible combinations of resource allocations to find the global optimal solution. If there are $|U|$ UEs, $|B|$ RUs, $M$ PRBs, and $P_{\text{levels}}$ number of discrete power levels, the total cost is given by:
Total $\text{Total Cost}=(|U|\times|B|)\times(M+1)!\times P_{\text{levels}}$
Where $(M+1)!$ represents the worst case of assigning UE to PRB or not, ESA needs to scale better with increasing network size and complexity due to its exponential growth in computational requirements. Despite its guarantee of finding the global optimal solution, it is impractical for real-time applications and large-scale networks.
VI-2 DQN
The complexity of training a DQN is expressed as $O(E\cdot(T\cdot(|S|\cdot|A|+F)))$, where $E$ is the number of episodes, $T$ is the number of steps per episode, $|S|$ is the size of the state space, $|A|$ is the size of the action space, and $F$ denotes the complexity of the neural network forward pass and backpropagation. The complexity of making a decision using a trained DQN is $O(F)$, where $F$ is the neural network forward pass complexity. The total computational cost for training and inference is dominated by the training phase, which can be expressed as:
Total $\text{Total Cost}=O(E\cdot(T\cdot(|S|\cdot|A|+F)))$
DQN can become computationally expensive as the state and action spaces grow due to the need for extensive exploration and training to converge to an optimal policy. This makes DQN less scalable for large, complex networks.
VI-3 SS-VAE
The complexity of training a VAE is $O(E\cdot(D\cdot L+L^{2}))$, where $E$ is the number of epochs, $D$ is the dimensionality of the input data, and $L$ is the number of latent dimensions. The complexity of inferring resource allocations using a trained VAE is $O(D\cdot L+L^{2})$. In addition, the complexity of training with contrastive loss is $O(N\cdot D\cdot L)$, where $N$ is the number of samples. Combining both phases, the total computational cost of the SS-VAE method is:
Total $\text{Total Cost}=O(E\cdot(D\cdot L+L^{2}))+O(N\cdot D\cdot L)$
The SS-VAE method scales efficiently with the number of features and samples due to its reliance on learning from data. While the training phase is computationally intensive, it is performed offline, and the inference phase is relatively lightweight, making it suitable for real-time applications in dynamic O-RAN environments.
VI-4 Comparison and Implications
The ESA guarantees finding the global optimal solution, but its computational infeasibility makes it impractical for large-scale networks. DQN provides a scalable solution; however, it requires extensive training and help with efficiency in large, complex networks. The SS-VAE method balances computational efficiency and scalability, making it suitable for real-time resource allocation in dynamic O-RAN environments. It achieves near-optimal performance with significantly lower computational costs, making it a robust and practical solution for modern O-RAN systems.
VII CONCLUSION
@@
This paper presents a novel model for the downlink of the O-RAN system, designed to coordinate multiple independent xAPPs to address the resource allocation problem with network slicing in the O-RAN architecture. It focuses on optimizing the aggregate weighted throughput by managing RU associations, power allocation, and PRBs to support two central 5G services, eMBB, and URLLC.
@@
Due to the non-convex and NP-hard nature of the resource allocation optimization problem, we proposed a semi-supervised learning approach to solve the resource allocation problem efficiently. Our algorithm begins by reformulating the problem as an optimization problem. The training phase involves supervised learning using a VAE to estimate power transmission and association indicators, followed by unsupervised learning using a contrastive loss approach to improve the model’s generalization and robustness.
We compared the SS-VAE method with benchmark algorithms ESA and DQN in various scenarios and discussed their computational complexity. The results indicate that our method performs similarly to ESA while outperforming DQN. We also examined scenarios for eMBB and URLLC services, considering their specific QoS requirements. Through simulations, we demonstrated the efficiency of our model, making it a robust candidate for resource allocation in O-RAN systems. Given the dynamic and complex nature of O-RAN environments, the sample efficiency of our model allows it to adapt to changing conditions without extensive retraining, ensuring high performance with fewer labeled data samples and enabling faster deployment and response times. The SS-VAE model provides robustness, reliability, and generalization from limited data.
@@
In our future research, we aim to explore several areas. Firstly, we intend to evaluate the end-to-end delay experienced by individual UEs and study the impact of deployment costs associated with network function placement. Addressing UE mobility is crucial in practical scenarios. Furthermore, we will explore alternative ML and optimization algorithms to improve our resource allocation approach, paving the way for new research opportunities in this dynamic field.